# Building sets, Chow rings, and their Hilbert series

### v1

@@ Abstract.
We establish formulas for the Hilbert series of the Feichtner–Yuzvinsky Chow ring of a polymatroid using arbitrary building sets. For braid matroids and minimal building sets, our results produce new formulas for the Poincaré polynomial of the moduli space ¯$\overline{\mathcal{M}}_{0,n+1}$ of pointed stable rational curves, and recover several previous results by Keel, Getzler, Manin, and Aluffi–Marcolli–Nascimento. We also use our methods to produce examples of matroids and building sets for which the corresponding Chow ring has Hilbert series with non-log-concave coefficients. This contrasts with the real-rootedness and log-concavity conjectures of Ferroni–Schröter for matroids with maximal building sets, and of Aluffi–Chen–Marcolli for braid matroids with minimal building sets.
@@ 1. Introduction
If furthermore $\operatorname{rk}(A)\leq|A|$ for every $A\subseteq E$, then $\mathsf{M}$ is a matroid. A flat of $\mathsf{M}$ is a subset $F\subseteq E$ such that $\operatorname{rk}(F\cup\{e\})>\operatorname{rk}(F)$ for all $e\in E\setminus F$. The set of all flats of $\mathsf{M}$ ordered by inclusion is a lattice, denoted $\mathcal{L}(\mathsf{M})$. We will always assume that a polymatroid $\mathsf{M}$ is loopless, i.e. that $\varnothing\in\mathcal{L}(\mathsf{M})$. We refer to [Wel76, Sch03] for detailed treatments of polymatroids.
A subset $\mathcal{G}\subseteq\mathcal{L}(\mathsf{M})$ of nonempty flats is a building set if it satisfies a factorability condition (see Definition 2.1). To any polymatroid $\mathsf{M}$ and building set $\mathcal{G}$, Feichtner and Yuzvinsky [FY04] associated a Chow ring $D(\mathsf{M},\mathcal{G})$ (see Definition 2.6), motivated by the geometry of wonderful compactifications of De Concini and Procesi [DCP95]. It is an Artinian graded ring $D(\mathsf{M},\mathcal{G})=\bigoplus_{i}D^{i}(\mathsf{M},\mathcal{G})$ whose Hilbert series is denoted
@@
  • When $\mathsf{M}$ is a matroid, the Chow ring $D(\mathsf{M},\mathcal{G}_{\max})$ with respect to the maximal building set $\mathcal{G}_{\max}$ is the Chow ring of the matroid $\mathsf{M}$ as given in [AHK18, Definition 1.3]. It is a central object in the Hodge theory of matroids [AHK18] that resolved the Heron–Rota–Welsh conjecture [Rot71, Her72, Wel76] on the log-concavity of the coefficients of the characteristic polynomial of a matroid.
  • When $\mathsf{M}$ is a braid matroid $\mathsf{K}_{n}$, i.e. the graphic matroid associated to a complete graph on $n$ vertices, the ring $D(\mathsf{K}_{n},\mathcal{G}_{\min})$ with respect to the minimal building set $\mathcal{G}_{\min}$ is isomorphic to the Chow ring of the Deligne–Knudsen–Mumford moduli space ¯$\overline{\mathcal{M}}_{0,n+1}$ of stable rational curves with $(n+1)$ marked points.
@@
  • For matroids with maximal building sets, see [BES24, BHM+22, FS24, FMSV24] for proofs of combinatorial interpretations, valuativity, and general recursions; [Stu24, FMV24] for inequalities for Chow polynomials in the broader context of graded posets; [ANR25, Lia24] for equivariant counterparts; and [HRS21, Hos24, BV25] for results that pertain to special cases such as uniform matroids.
  • For the braid matroid $\mathsf{K}_{n}$, three different formulas for the Hilbert series $\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ have been found by Manin [Man95], Keel [Kee92], and more recently by Aluffi, Marcolli, and Nascimento [AMN24], all via the geometry of the space ¯$\overline{\mathcal{M}}_{0,n+1}$.
The ring $D(\mathsf{M},\mathcal{G})$ is known to satisfy a trio of properties known as the Kähler package [PP23, Section 4]; see also [ADH23] and [CHL+25]. One of the components of the package, the Hard Lefschetz theorem, implies that the sequence of coefficients of the polynomial $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ is symmetric and unimodal. Under the additional assumption that $\mathsf{M}$ is a matroid and $\mathcal{G}$ is the maximal building set, it was conjectured by Ferroni and Schröter in [FS24] that the corresponding Hilbert series are real-rooted polynomials. Real-rootedness is the strongest in the hierarchy of properties depicted in Figure 1 (see [Brä15, Bre94, Sta89]). The weaker property of $\gamma$-positivity was established in this special case [FMSV24].
@@
In the general setting of arbitrary building sets, all of the properties appearing in Figure 1, except unimodality, fail to be true. The failure of real-rootedness and $\gamma$-positivity is easily found: when $\mathsf{M}$ is a uniform matroid of rank $k\geq 3$ and $\mathcal{G}=\mathcal{G}_{\min}$, one has $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=1+x+\cdots+x^{k-1}$. The failure of log-concavity will be given in Theorem 1.6 (with more details in Section 6). The construction of this example is an application of some recursions we develop that generalize those in [FMSV24] to the case of arbitrary building sets.
@@ 1.1. Main results
where $\mu$ denotes the Möbius function of the poset $\mathcal{L}(\mathsf{M})$ (see [Sta12, Chapter 3]). The set of $\mathcal{G}$-factors of a building set $\mathcal{G}$ on $\mathsf{M}$, denoted $f(\mathcal{G})$, consists of all maximal elements in the subset $\mathcal{G}$ of the poset $\mathcal{L}(\mathsf{M})$. For a flat $F$ of $\mathsf{M}$, a building set $\mathcal{G}$ naturally induces building sets $\mathcal{G}|_{F}$ and $\mathcal{G}/F$ on the restriction $\mathsf{M}|_{F}$ and the contraction $\mathsf{M}/F$ polymatroids, respectively (see Definition 2.3). If $\mathsf{M}$ is the empty polymatroid, we set $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=1$.
@@
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{\varnothing\neq F\in\mathcal{L}(\mathsf{M})}\frac{-\chi_{\mathsf{M}|_{F}}(x)}{(1-x)^{|f(\mathcal{G}|_{F})|}}\cdot\mathrm{H}_{\mathsf{M}/F}^{\mathcal{G}/F}(x),$ (1)
@@
$\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\mathrm{H}_{\mathsf{M}|_{F}}^{\mathcal{G}|_{F}}(x)\cdot\frac{-\chi_{\mathsf{M}/F}(x)}{(1-x)^{\lvert f(\mathcal{G}/F)\rvert}}.$ (2)
We prove these recursions using the main result of Feichtner and Yuzvinsky in [FY04], which consists of a Gröbner basis computation leading to a basis of $D(\mathsf{M},\mathcal{G})$ as a $\mathbb{Q}$-vector space. Theorem 1.1 is equivalent to the statement (Theorem 3.6) that two particular elements in the incidence algebra of $\mathcal{L}(\mathsf{M})$ are inverses. Thus, using the formalism of incidence algebras, we derive a non-recursive formula as a sum over chains of flats.
@@
The formulas in Corollary 1.2 and Theorem 1.3 are useful for computing Hilbert series of rings $D(\mathsf{M},\mathcal{G})$ when the characteristic polynomials of the intervals of $\mathcal{L}(\mathsf{M})$ obey a predictable pattern. A central example is when $\mathsf{M}=\mathsf{K}_{n}$: in this case, the reduced characteristic polynomials of intervals of flats are products of polynomials of the form $(x-2)(x-3)\cdots(x-r)$. When $\mathcal{G}=\mathcal{G}_{\min}$ is the minimal building set, the wonderful variety (see Section 3.5) is the moduli space ¯$\overline{\mathcal{M}}_{0,n+1}$. Our results lead us to a new closed formula for its Poincaré polynomial.
@@
The Poincaré polynomial of ¯$\overline{\mathcal{M}}_{0,n+1}$ is
@@
where the sum is over set partitions of the set $[n-1]=\{1,\dotsc,n-1\}$ into $\ell(\lambda)$ nonempty parts, $\lambda_{i}$ is the cardinality of the $i$-th part, and $\overline{\chi}_{m}(x):=(x-2)(x-3)\dotsb(x-m+1)$.
@@
Theorem 1.5 ([AMN24, Theorem 1.1])
The Poincaré polynomial of ¯$\overline{\mathcal{M}}_{0,n+1}$ satisfies
@@
For the Chow ring of a matroid $\mathsf{M}$ with the maximal building set $\mathcal{G}_{\max}$, an open conjecture of Ferroni–Schröter [FS24] posits that the Hilbert series $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}_{\max}}(x)$ is real-rooted. However, for general building sets this property and the weaker property of $\gamma$-positivity do not hold. We provide a negative answer to the analogous question for log-concavity.
@@ Acknowledgments
We are grateful to the organizers of the Oberwolfach workshop “Arrangements, matroids, and logarithmic vector fields”—this project was initiated at that MFO workshop. The authors also thank Emil Verkama for insightful discussions. Chris Eur was partially supported by the National Science Foundation grant DMS-2246518. Luis Ferroni is a member at the Institute for Advanced Study, funded by the Minerva Research Foundation. Jacob Matherne received support from a Simons Foundation Travel Support for Mathematicians Award MPS-TSM-00007970. Roberto Pagaria is partially supported by PRIN 2022 “ALgebraic and TOPological combinatorics (ALTOP)” CUP J53D23003660006 and by INdAM - GNSAGA Project CUP E53C23001670001.
@@ 2.1. Building sets
$f(\mathcal{G}):=\text{the set of maximal elements in the subset $\mathcal{G}$ of the poset $\mathcal{L}(\mathsf{M})$}.$
@@
Let $\mathsf{M}$ be a loopless polymatroid, and let $\mathcal{L}(\mathsf{M})$ be its lattice of flats. A subset $\mathcal{G}\subseteq\mathcal{L}(\mathsf{M})\setminus\{\varnothing\}$ is said to be a (geometric) building set if for any $F\in\mathcal{L}(\mathsf{M})$ the morphism of lattices
@@
We emphasize that we do not require the ground set $E$ to be part of a building set, in contrast to some other works. Our definition of a building set coincides with the original definition by De Concini and Procesi [DCP95].
@@
For a flat $F$ of a polymatroid $\mathsf{M}=(E,\operatorname{rk})$, the restriction $\mathsf{M}|_{F}$ is the polymatroid $(F,\operatorname{rk}|_{F})$, and the contraction $\mathsf{M}/F$ is the polymatroid $(E\setminus F,\operatorname{rk}^{\prime})$ where $\operatorname{rk}^{\prime}(A):=\operatorname{rk}(A\cup F)-\operatorname{rk}(F)$. The lattices of flats of $\mathsf{M}|_{F}$ and $\mathsf{M}/F$ are isomorphic to the intervals $[\varnothing,F]$ and $[F,E]$ of $\mathcal{L}(\mathsf{M})$, respectively. These operations extend to building sets, as follows.
@@
If $\mathcal{G}=\mathcal{G}_{\max}$ is the maximal building set for $\mathsf{M}$, then $\mathcal{G}|_{F}$ and $\mathcal{G}/F$ are maximal building sets of the restriction and contraction respectively. We caution however that when $\mathcal{G}=\mathcal{G}_{\min}$ is the minimal building set, $\mathcal{G}/F$ may not be the minimal building set in the contraction, since contracting a flat may not preserve connectivity.
@@ 2.2. Nested sets and Chow rings
The central objects of study in this paper are the Chow rings arising from a polymatroid and a building set. These rings (in a slightly more general set up) were introduced by Feichtner and Yuzvinsky in [FY04].
@@
where $I$ is the homogeneous ideal generated by the relations
[…]
and
$\sum_{G\geq F}x_{G}\qquad\text{for $F$ an atom of $\mathcal{L}(\mathsf{M})$}.$
@@
One of the main results by Feichtner and Yuzvinsky, which follows from the construction of a Gröbner basis for the ideal $I$ in Definition 2.6, is that the rings $D(\mathsf{M},\mathcal{G})$ possess an additive basis given by a specific list of monomials. Since this is useful for computing the Hilbert series, we restate it here.
Theorem 2.7 ([FY04, Corollary 1] and [PP23, Corollary 2.8])
@@
The monomials appearing in the preceding statement will be customarily called the “FY-monomials” associated to $\mathsf{M}$ and $\mathcal{G}$. Notice that when $\mathcal{G}=\mathcal{G}_{\max}$, a nested set is just a flag of flats $S=\{F_{1}\subsetneq\cdots\subsetneq F_{m}\}$, and the rank condition in the theorem reads $10\leq e_{i}\leq\operatorname{rk}(F_{i})-\operatorname{rk}(F_{i-1})-$ 1 for all $1\leq i\leq m$.

### v2

@@ Abstract.
We establish formulas for the Hilbert series of the Chow ring of a polymatroid using arbitrary building sets. For braid matroids and minimal building sets, our results produce new formulas for the Poincaré polynomial of the moduli space $\overline{\mathcal{M}}_{0,n+1}$ of pointed stable rational curves, and recover several previous results by Keel, Getzler, Manin, and Aluffi–Marcolli–Nascimento. We also use our methods to produce examples of matroids and building sets for which the corresponding Chow ring has Hilbert series with non-log-concave coefficients. This contrasts with the real-rootedness and log-concavity conjectures of Ferroni–Schröter for matroids with maximal building sets, and of Aluffi–Chen–Marcolli for braid matroids with minimal building sets.
@@ 1. Introduction
If furthermore $\operatorname{rk}(A)\leq|A|$ for every $A\subseteq E$, then $\mathsf{M}$ is a matroid. A flat of $\mathsf{M}$ is a subset $F\subseteq E$ such that $\operatorname{rk}(F\cup\{e\})>\operatorname{rk}(F)$ for all $e\in E\setminus F$. The set of all flats of $\mathsf{M}$ ordered by inclusion is a lattice, denoted $\mathcal{L}(\mathsf{M})$. We will always assume that a polymatroid $\mathsf{M}$ is loopless, i.e. that $\varnothing\in\mathcal{L}(\mathsf{M})$. We refer to [44, 39] for detailed treatments of polymatroids.
A subset $\mathcal{G}\subseteq\mathcal{L}(\mathsf{M})$ of nonempty flats is a building set if it satisfies a factorability condition (see Definition 2.1). To any matroid $\mathsf{M}$ and building set $\mathcal{G}$, Feichtner and Yuzvinsky [20] associated a Chow ring $D(\mathsf{M},\mathcal{G})$ (see Definition 2.6), motivated by the geometry of wonderful compactifications of De Concini and Procesi [15]; this construction was later extended to polymatroids with arbitrary building sets by Pagaria and Pezzoli in [36]. These are Artinian graded rings $D(\mathsf{M},\mathcal{G})=\bigoplus_{i}D^{i}(\mathsf{M},\mathcal{G})$ whose Hilbert series are denoted
@@
  • When $\mathsf{M}$ is a matroid, the Chow ring $D(\mathsf{M},\mathcal{G}_{\max})$ with respect to the maximal building set $\mathcal{G}_{\max}$ is the Chow ring of the matroid $\mathsf{M}$ as given in [1, Definition 1.3]. It is a central object in the Hodge theory of matroids [1] that resolved the Heron–Rota–Welsh conjecture [38, 28, 44] on the log-concavity of the coefficients of the characteristic polynomial of a matroid.
  • When $\mathsf{M}$ is a braid matroid $\mathsf{K}_{n}$, i.e. the graphic matroid associated to a complete graph on $n$ vertices, the ring $D(\mathsf{K}_{n},\mathcal{G}_{\min})$ with respect to the minimal building set $\mathcal{G}_{\min}$ is isomorphic to the Chow ring of the Deligne–Knudsen–Mumford moduli space $\overline{\mathcal{M}}_{0,n+1}$ of stable rational curves with $(n+1)$ marked points.
@@
  • For matroids with maximal building sets, see [6, 7, 23, 21] for proofs of combinatorial interpretations, valuativity, and general recursions; [43, 22] for inequalities for Chow polynomials in the broader context of graded posets; [4, 32] for equivariant counterparts; and [27, 29, 8] for results that pertain to special cases such as uniform matroids.
  • For the braid matroid $\mathsf{K}_{n}$, three different formulas for the Hilbert series $\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ have been found by Manin [33], Keel [31], and more recently by Aluffi, Marcolli, and Nascimento [3], all via the geometry of the space $\overline{\mathcal{M}}_{0,n+1}$.
The ring $D(\mathsf{M},\mathcal{G})$ is known to satisfy a trio of properties known as the Kähler package [36, Section 4]; see also [5] and [14]. One of the components of the package, the Hard Lefschetz theorem, implies that the sequence of coefficients of the polynomial $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)$ is symmetric and unimodal. Under the additional assumption that $\mathsf{M}$ is a matroid and $\mathcal{G}$ is the maximal building set, it was conjectured by Ferroni and Schröter in [23] that the corresponding Hilbert series are real-rooted polynomials. Real-rootedness is the strongest in the hierarchy of properties depicted in Figure 1 (see [9, 11, 40]). The weaker property of $\gamma$-positivity was established in this special case [21].
@@
In the general setting of arbitrary building sets, all of the properties appearing in Figure 1, except unimodality, fail to be true. The failure of real-rootedness and $\gamma$-positivity is easily found: when $\mathsf{M}$ is a uniform matroid of rank $k\geq 3$ and $\mathcal{G}=\mathcal{G}_{\min}$, one has $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=1+x+\cdots+x^{k-1}$. The failure of log-concavity will be given in Theorem 1.6 (with more details in Section 6). The construction of this example is an application of some recursions we develop that generalize those in [21] to the case of arbitrary building sets.
@@ 1.1. Main results
where $\mu$ denotes the Möbius function of the poset $\mathcal{L}(\mathsf{M})$ (see [42, Chapter 3]). In the case of a matroid realized as a complex hyperplane arrangement, the characteristic polynomial of a matroid has, up to an alternating factor, the same coefficients as the Poincaré polynomial of the cohomology of the arrangement complement (see [34]). The set of $\mathcal{G}$-factors of a building set $\mathcal{G}$ on $\mathsf{M}$, denoted $f(\mathcal{G})$, consists of all maximal elements in the subset $\mathcal{G}$ of the poset $\mathcal{L}(\mathsf{M})$. For a flat $F$ of $\mathsf{M}$, a building set $\mathcal{G}$ naturally induces building sets $\mathcal{G}|_{F}$ and $\mathcal{G}/F$ on the restriction $\mathsf{M}|_{F}$ and the contraction $\mathsf{M}/F$ polymatroids, respectively (see Definition 2.3). If $\mathsf{M}$ is the empty polymatroid, we set $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=1$.
@@
(1) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{\varnothing\neq F\in\mathcal{L}(\mathsf{M})}\frac{-\chi_{\mathsf{M}|_{F}}(x)}{(1-x)^{|f(\mathcal{G}|_{F})|}}\cdot\mathrm{H}_{\mathsf{M}/F}^{\mathcal{G}/F}(x),$
@@
(2) $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}}(x)=\sum_{E\neq F\in\mathcal{L}(\mathsf{M})}\mathrm{H}_{\mathsf{M}|_{F}}^{\mathcal{G}|_{F}}(x)\cdot\frac{-\chi_{\mathsf{M}/F}(x)}{(1-x)^{\lvert f(\mathcal{G}/F)\rvert}}$.
We prove these recursions using a result of Pagaria–Pezzoli in [36] which consists of a Gröbner basis computation leading to a basis of $D(\mathsf{M},\mathcal{G})$ as a $\mathbb{Q}$-vector space. Theorem 1.1 is equivalent to the statement (Theorem 3.6) that two particular elements in the incidence algebra of $\mathcal{L}(\mathsf{M})$ are inverses. Thus, using the formalism of incidence algebras, we derive a non-recursive formula as a sum over chains of flats.
@@
In the above theorem $\overline{\chi}_{\mathsf{M}}(x)$ stands for the reduced characteristic polynomial of $\mathsf{M}$, which is defined as $\frac{1}{x-1}\chi_{\mathsf{M}}(x)$ for each non-empty polymatroid $\mathsf{M}$. The formulas in Corollary 1.2 and Theorem 1.3 are useful for computing Hilbert series of rings $D(\mathsf{M},\mathcal{G})$ when the characteristic polynomials of the intervals of $\mathcal{L}(\mathsf{M})$ obey a predictable pattern. A central example is when $\mathsf{M}=\mathsf{K}_{n}$: in this case, the reduced characteristic polynomials of intervals of flats are products of polynomials of the form $(x-2)(x-3)\cdots(x-r)$. When $\mathcal{G}=\mathcal{G}_{\min}$ is the minimal building set, the wonderful variety (see Section 3.5) is the moduli space $\overline{\mathcal{M}}_{0,n+1}$. Our results lead us to a new closed formula for its Poincaré polynomial.
@@
The Poincaré polynomial of $\overline{\mathcal{M}}_{0,n+1}$ is
@@
where the sum is over set partitions of the set $n-1]=\{1,\dotsc,n-1\}$ into $\ell(\lambda)$ nonempty parts, $\lambda_{i}$ is the cardinality of the $i$-th part, and $\overline{\chi}_{m}(x):=(x-2)(x-3)\dotsb(x-m+1)$.
@@
Theorem 1.5 ([3, Theorem 1.1])
The Poincaré polynomial of $\overline{\mathcal{M}}_{0,n+1}$ satisfies
@@
For the Chow ring of a matroid $\mathsf{M}$ with the maximal building set $\mathcal{G}_{\max}$, an open conjecture of Ferroni–Schröter [23] posits that the Hilbert series $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}_{\max}}(x)$ is real-rooted. However, for general building sets this property and the weaker property of $\gamma$-positivity do not hold. We provide a negative answer to the analogous question for log-concavity.
@@ Acknowledgments
We are grateful to the organizers of the Oberwolfach workshop “Arrangements, matroids, and logarithmic vector fields”—this project was initiated at that MFO workshop. The authors also thank Emil Verkama for insightful discussions. We are also grateful to the anonymous referee for many helpful comments and suggestions. Chris Eur was partially supported by the National Science Foundation grant DMS-2246518. Luis Ferroni was a member at the Institute for Advanced Study, funded by the Minerva Research Foundation. Jacob Matherne received support from NSF Grant DMS-2452179 and Simons Foundation Travel Support for Mathematicians Award MPS-TSM-00007970. Roberto Pagaria is partially supported by PRIN 2022 “ALgebraic and TOPological combinatorics (ALTOP)” CUP J53D23003660006 and by INdAM - GNSAGA Project CUP E53C23001670001.
@@ 2.1. Building sets
[…]
@@
Let $\mathsf{M}$ be a loopless polymatroid, and let $\mathcal{L}(\mathsf{M})$ be its lattice of flats. A subset $\mathcal{G}\subseteq\mathcal{L}(\mathsf{M})\setminus\{\varnothing\}$ is said to be a building set if for any $F\in\mathcal{L}(\mathsf{M})$ the morphism of lattices
@@
We emphasize that we do not require the ground set $E$ to be part of a building set, in contrast to some other works. Our definition of a building set coincides with the original definition by De Concini and Procesi [15].
@@
For a flat $F$ of a polymatroid $\mathsf{M}=(E,\operatorname{rk})$, the restriction $\mathsf{M}|_{F}$ is the polymatroid $(F,\operatorname{rk}|_{F})$, and the contraction $\mathsf{M}/F$ is the polymatroid $(E\setminus F,\operatorname{rk}^{\prime})$ where $\operatorname{rk}^{\prime}(A):=\operatorname{rk}(A\cup F)-\operatorname{rk}(F)$. The lattices of flats of $\mathsf{M}|_{F}$ and $\mathsf{M}/F$ are isomorphic to the intervals $\varnothing,F]$ and $[F,E]$ of $\mathcal{L}(\mathsf{M})$, respectively. These operations extend to building sets, as follows.
@@
If $\mathcal{G}=\mathcal{G}_{\max}$ is the maximal building set for $\mathsf{M}$, then $\mathcal{G}|_{F}$ and $\mathcal{G}/F$ are maximal building sets of the restriction and contraction respectively. We caution, however, that when $\mathcal{G}=\mathcal{G}_{\min}$ is the minimal building set, $\mathcal{G}/F$ may not be the minimal building set in the contraction, since contracting a flat may not preserve connectivity.
@@ 2.2. Nested sets and Chow rings
The central objects of study in this paper are the Chow rings arising from a polymatroid and a building set. These rings (in a slightly more general set up) were introduced by Feichtner and Yuzvinsky in [20], and later generalized to polymatroids by Pagaria and Pezzoli [36]. The following definition can be found in [36, Section 4].
@@
where $I$ is the ideal generated by the relations
$\left(\prod_{i=1}^{m}x_{G_{i}}\right)\left(\sum_{\begin{subarray}{c}H\in\mathcal{G}\\ H\geq G\end{subarray}}x_{H}\right)^{b}\quad\text{for each subset $S=\{G_{1},\ldots,G_{m}\}\subseteq\mathcal{G}$ and $G\in\mathcal{G}$,}$
where $b:=\operatorname{rk}(G)-\operatorname{rk}(\sup(S,G))$ and $\sup(S,G):=\displaystyle\bigvee_{\begin{subarray}{c}G^{\prime}\in S\\ G^{\prime}<G\end{subarray}}G^{\prime}$.
@@
One of the main results by Pagaria–Pezzoli constructs a Gröbner basis for the ideal $I$ in Definition 2.6. In particular, the rings $D(\mathsf{M},\mathcal{G})$ possess an additive basis given by a specific list of monomials. Since this is useful for computing the Hilbert series, we restate it here.
Theorem 2.7 ([36, Corollary 2.8])
@@
The monomials appearing in the preceding statement will be customarily called the “FY-monomials” associated to $\mathsf{M}$ and $\mathcal{G}$. Notice that when $\mathcal{G}=\mathcal{G}_{\max}$, a nested set is just a flag of flats $S=\{F_{1}\subsetneq\cdots\subsetneq F_{m}\}$, and the rank condition in the theorem reads $0\leq e_{i}\leq\operatorname{rk}(F_{i})-\operatorname{rk}(F_{i-1})-1$ for all $1\leq i\leq m$.